# Warming-driven rise in soil moisture entropy signals destabilization of the Asian Water Tower

### v1

@@ 4.4 Local Facilitation Cellular Automaton model (LFCA)
Data availability The GLDAS-NOAH data used here are publicly available at https://disc.gsfc.nasa.gov/datasets/GLDAS_NOAH025_M_2.1/summary?keywords=GLDAS. The ERA5 reanalysis data are publicly available at https://cds.climate.copernicus.eu/datasets/reanalysis-era5-pressure-levels?tab=download. The MERRA-2 reanalysis data are publicly available at https://disc.gsfc.nasa.gov/datasets/M2TMNXLND_5.12.4/summary?keywords=M2CONXLND. The ERA5-Land reanalysis data are publicly available at https://cds.climate.copernicus.eu/datasets/reanalysis-era5-land-monthly-means?tab=download. The CMIP6 data are publicly available at https://aims2.llnl.gov/search. The ONI data are publicly available at https://origin.cpc.ncep.noaa.gov/products/analysis_monitoring/ensostuff/ONI_v5.php The MEI.v2 data are publicly available at https://psl.noaa.gov/enso/mei/.

### v2

@@ 4.4 Local Facilitation Cellular Automaton model (LFCA)
Data availability The GLDAS-NOAH data used here are publicly available at https://disc.gsfc.nasa.gov/datasets/GLDAS_NOAH025_M_2.1/summary?keywords=GLDAS. The ERA5 reanalysis data are publicly available at https://cds.climate.copernicus.eu/datasets/reanalysis-era5-pressure-levels?tab=download. The MERRA-2 reanalysis data are publicly available at https://disc.gsfc.nasa.gov/datasets/M2TMNXLND_5.12.4/summary?keywords=M2CONXLND. The ERA5-Land reanalysis data are publicly available at https://cds.climate.copernicus.eu/datasets/reanalysis-era5-land-monthly-means?tab=download. The CMIP6 data are publicly available at https://aims2.llnl.gov/search. The ONI data are publicly available at https://origin.cpc.ncep.noaa.gov/products/analysis_monitoring/ensostuff/ONI_v5.php. The MEI.v2 data are publicly available at https://psl.noaa.gov/enso/mei/.